### v1

@@ 4.2.2. Qualitative Results
Workshop data presented debates centering on delineating legal information versus advice. Most participants deemed offering factual legal information acceptable, but some participants envisioned using the conversational strengths of LLMs to help users clarify inquiries and pinpoint applicable laws through follow-up questioning. Furthermore, experts suggested additional layers of ethical guidelines regarding appropriate AI responses.
@@
  • Do not make factual determinations. Providing relevant laws is fine (e.g., driving under influence (DUI) is illegal in Washington) but applying it to specific user situations constitutes opinion (e.g., falling asleep in the driver’s seat in the parking lot after drinking alcohol could be a DUI) (P2, P13, P17, P19).
@@
In essence, legal opinion encompasses interpretive, judgment-driven analysis that is often value-laden and forward-looking, whereas legal information primarily involves reporting objective laws and past rulings without subjective assessment. The distinction can be informed by legal tradition, such as the IRAC (Issue-Rule-Analysis-Conclusion) framework. Despite its mysterious origin, IRAC is the most ubiquitous legal reasoning method taught to law students and employed by lawyers drafting memos (Metzler, 2002). At its core, IRAC is the application of deductive logic to the law. It entails identifying the salient legal issue, stating the governing rule that applies, analyzing how the particular facts interact with the stipulations of the rule, and finally deducing the inevitable conclusion (Center, 2001). Our findings show that AI systems focusing on issue and rule identification provide permissible fact-finding “information,” whereas analysis and conclusions may cross into tailored “opinion,” as shown in Figure 5. This demonstrates how principles accumulated over decades of legal practice inform AI policy.
@@
While cautioning against detailed legal opinions, participants indicated AI could still improve on search engines within ethical boundaries. As P20 noted, users would not welcome AI systems’ “vomiting a whole lot of knowledge.” The most promising and heavily-discussed area is enabling multi-turn interactions, or AI systems asking follow-up questions to clarify users’ legally meaningful questions. This possibility emerged as participants expressed frustration on missing case facts: “I don’t think there’s enough information to go off of, and that depending on the details that come out, it could change the analysis, therefore the outcomes.” (P13) Because contexts are always complex (P11), lawyers take time to elicit relevant facts and identify the “right questions to ask” (P12). Participants felt such dialogues could streamline “screening interviews” (P12), “first calls” (P14), or “intake meetings” (P15) as opposed to one-shot searches. This process could help users to focus on key aspects and connect them to relevant expertise.
However, some warned developers should exercise care before eliciting extensive personal information (P13, P16), given confidentiality concerns. While issue-spotting and rule identification likely constitute permissible legal information, the line between information and advice remains blurred. P16 viewed that narrowing down factual patterns and rules engages deep judgement; “you’re starting to make the AI become your lawyer.” This claim is supported by the current legal uncertainty of unauthorized practice of law (UPL), where certain states construe the law stringently, even restricting self-explanatory informational materials (boo, 1967). Therefore, while iterative legal clarification offers value, AI systems may decide to err conservatively rather than overreach.
@@
Experts suggested several principles for responsibly providing AI legal assistance. Some principles directly align with emerging literature on transparency (Kim et al., 2023a), user satisfaction (Kim et al., 2023b), and cautions about anthropomorphism (Shanahan et al., 2023; Ma et al., 2023). Please note that the following list does not reflect full consensus across interviews.
  • Don’t Pretend to Be Human: AI systems should not behave like a human and cause misrepresentations, as that can create issues around transparency, over-reliance, and managing user expectations.
  • Caveat Constraints: AI systems should provide various caveats on its limitations, such as that its capabilities are constrained, the conversation is not privileged, and it is working off of incomplete information.
  • Avoid Potential Harm: AI systems should avoid recommendations that could potentially cause harm, both in terms of real-world actions a user takes based on guidance, but also emotional or psychological harm from its own responses.
  • Respect the Justice System: AI systems should not give answers that could enable users to intentionally avoid law enforcement or oversight.
  • Avoid Unethical Answers: AI systems should not make any outputs that could promote dishonesty or deception, fraud, impersonation, or other unethical behaviors that could get users into legal trouble.
  • Be Transparent: AI systems should be able to explain the outcome it generated and point to the specific areas of code or data it relied on.
  • Avoid Appearance of Impropriety: AI systems should avoid the appearance of conflicts of interest such as endorsing its creators or AI companies in general.
@@ 4.3. Summary of Results
Our analysis uncovered 25 distinct dimensions pertinent to ensuring safe and effective AI legal assistance, spanning four key categories: (1) user attributes and behaviors, (2) query characteristics, (3) current AI capabilities, and (4) impact considerations. Experts deliberated with each other and through points of consensus to produce this rich set of considerations. However, experts expressed limited consensus on how AI systems should actually respond, given these nuanced factors. Some remained resistant to any AI involvement in legal questions, while others envisioned more helpful AI assistance models that increase access to information. Most debates surrounded distinguishing information versus opinion, and the majority felt that providing factual legal information is appropriate. Some participants envisioned using LLMs’ conversational capabilities to help users refine questions and identify relevant laws through follow-up questions, similar to initial consultations with attorneys.
@@ 5. Discussion
Constructing AI policy does not exist in a technocratic silo—instead, responsible AI legal advice requires cross-disciplinary synthesis involving domain experts. We demonstrate that engaging legal experts in case-based deliberation can successfully translate professional knowledge and clinical experience into a concrete set of considerations for AI policy. Our 4-dimension framework provides an analytical framework suitable for further research in the legal domain as well as other professional domains. Overall, we argue that centuries-old legal traditions prove useful in informing AI policy in the contemporary age.
Benefits of Case-based Deliberation Methods
@@
Our research process underscores several advantages of grounded case deliberation for eliciting considerations. Preparing realistic scenarios, while laborious, proved invaluable in quickly engaging experts with targeted queries related to their clinical experience. The realistic cases allowed experts to examine granular concerns around singular situations as well as overarching technical and legal constraints, producing a more concrete set of contextual factors for AI developers, beyond theoretical and high-level principles in prior research (Solaiman et al., 2023; Antoniak et al., 2023; Ganguli et al., 2022). Finally, the collective deliberation itself revealed critical hidden dimensions and elicited justifications that shed new light on existing dimensions. As experts built on each other’s points, they realized overlooked issues or limitations in their own initial analyses. This interplay sharpened considerations and underscored nuances around balancing risks and benefits in varied situations. The combination of grounded cases and collaborative discourse resulted in more fine-grained, practice-informed insights compared to de-contextualized surveys or high-level principles.
Applicability to Other Professional Domains
While each possessing unique dimensions, domains like medicine, mental health, law, and finance share common threads around high-stakes real-world impact and historical reliance on licensed specialists for advice. We believe that our research methods and 4-dimension framework give illustrative guidance to further research in other professional domains. As this research demonstrates how case-based elicitation methods can unravel complex professional ethics, researchers could adopt similar processes engaging mental health counselors, financial advisors, or medical professionals. Tapping into the clinical experience and ethical knowledge of practitioners through structured deliberation based on realistic cases can help produce tailored dimensions and guidelines for responsible AI assistance respective to each profession.
Furthermore, our 4-dimension framework—(1) user, (2) query, (3) AI capabilities, and (4) impact—could be used across professional domains. Considerations within the user, AI, and impact categories, which resembles human-AI-context categories in AI trustworthiness literature (Kim et al., 2023a), can also be applied in other domains with minimal changes. However, the query dimensions warrants more customization to address process and outcome characteristics in specific fields. For example, the applicable regulations, typical requests, terminologies, and satisfactory responses in personal investing diverge from constructing legal briefs, medical care plans, or emotional counseling. While some dimension themes persist, their precise formulation specifics requires reconceptualization by experts in domain-specific contexts.
Charting Novel Legal Considerations
Despite rapid advances, AI systems currently cannot replace human legal counsel. One of our contributions is to shed light on existing legal and ethical barriers to AI’s legal advice which have been overlooked in LLM literature. For one, Section 4.1.3 reveals that users lack confidentiality and accountability protections governing attorney advice—conversations with AI systems risk disclosure in legal proceedings and inaccurate guidance evades professional negligence liability. Moreover, as Section 4.2.2 explains, UPL regulations prohibit non-lawyers from advising in many states, carrying criminal penalties. In other words, uncontrolled rollout of AI guidance may constitute an illegal practice with harsh consequences.
However, legal conservatism could evolve as AI systems advance in reasoning and precision (Trautmann et al., 2022; Jiang and Yang, 2023; Louis et al., 2023; Prasad et al., 2022; Xiao et al., 2021; Bakker et al., 2022; Yue et al., 2023). For instance, UPL rules have faced past criticism for limiting affordable access to legal help (Denckla, 1998; Rotenberg, 2012; on Nonlawyer Practice, 1995), and the EU AI Act categorizes AI legal assistance tools as “high-risk” which subjects them to heightened responsibilities instead of banning them (European Commission, 2021). In the face of legal changes, tracking emerging use cases and legal boundaries prove critical not only for regulatory compliance but also for more responsible innovation. We could imagine AI systems designed like private counsels advising single parties, rather than serving all users uniformly like ChatGPT. In such case, AI could come to resemble proprietary services, with corresponding confidentiality and liability assurances.
Learning from Professional Communities
@@
As emerging research on NLP in healthcare emphasizes (Antoniak et al., 2023), learning from communities with centuries of expertise can help sidestep painful mistakes. For instance, UPL does not only constrain the possibility of AI’s legal advice but also proposes a time-tested distinction criterion between information versus opinion. Merely providing legal information has not been historically punished as a UPL violation (cal, 1970; Denckla, 1998; cal, 2007). For example, the Texas Court provides guidelines for court staff and illustrative examples like in Table 2. These examples show subtle difference between information and opinion, which resembles the red-teaming approach to distinguish harmful user prompts (Ganguli et al., 2022; Achiam et al., 2023). Furthermore, legal scholars have explored legally justifiable AI advice under UPL, attorney-client privilege, and other doctrines (Haupt, 2019; Wendel, 2019; Spahn, 2017). Wendel’s definition of the “core lawyering functions” is a pertinent one: Important tasks like recommending the course of actions or drafting contracts cannot be delegated to AI agents due to technical limitations and accountability deficits (Wendel, 2019). This demonstrates how principles accumulated over centuries of legal practice now inform responsible AI systems and the call for cross-disciplinary collaborations.
@@
Procedure
Can you tell me how to file a small claims action?
Can you tell me whether it would be better to file a small claims action or a civil action?
Definition
What does “certificate of service” mean?
My neighbors leave their kids at home all day without supervision. Isn’t that child neglect?
Forms
I need to file for divorce and I have no idea where to begin. Is there some place I can go to find out how to get started?
The self-help divorce petition says I should list any gifts as my separate property. Should I list the money that my parents gave me last month as my separate property?
Options
What can I do if I cannot afford to pay the filing fee?
My ex-husband hasn’t paid the debts that he agreed to pay in our divorce settlement. Can I be made responsible for this debt?
Table 2. Examples of impermissible questions that requires legal opinion (tex, 2015). Remarkably similar to how red-teaming in LLM development identifies harmful user inputs (Ganguli et al., 2022; Achiam et al., 2023), this edited list (compiled from Texas law clerk resources) distinguishes between permissible and forbidden questions Texas court personnel can answer.
Limitations and Future Research
Our study has several limitations. First, our expert sample predominantly focused on practitioners familiar with the US legal system. Ethical considerations around appropriate AI assistance may differ significantly across legal systems and cultures around the world. Second, we did not directly engage end-users like clients of legal services. Future work can specifically investigate end-user perceptions to compare and contrast with our expert-informed results. Finally, while our taxonomy conceptualizes a concrete set of dimensions, how these dimensions could change the appropriateness of AI responses remains unexplained. This may require larger-scale empirical analysis on public assessments across diverse pairings of cases and responses.
@@ 6. Conclusion
Today, LLM chatbots are increasingly capable of providing users with advice in a wide range of professional domains, including legal advice. However, what constitutes an appropriate AI-generated response to legal queries, where both required expertise and resulting consequences are high? To explore this, we conducted workshops with 20 legal experts using methods inspired by case-based reasoning to encourage deliberations around appropriate AI responses to legal queries in practice. Our contributions are threefold. First, we presented a set of 25 key dimensions, synthesized from expert deliberations, that impacted AI response appropriateness in the legal domain. Second, we shared experts’ recommendations for AI response strategies and guiding principles for generating appropriate responses—these centered around helping users identify and prepare salient information for legal proceedings rather than to recommend specific legal actions. Finally, we posit that our case-based method has utility in engaging expert perspectives on AI response appropriateness in professional domains beyond the legal sphere. Taken together, our work sets an empirical foundation for translating domain-specific professional knowledge and practices into AI policies to steer real-world AI behavior in a more responsible direction.

### v2

@@ 4.2.2. Qualitative Results
Our qualitative analysis revealed rich and nuanced discussions behind the categorization of desired LLM responses. Experts delved into the complexities of distinguishing between legal information and opinion and the challenges of ensuring user protection while leveraging the capabilities of LLMs. They emphasized the importance of transparency, user safeguards, and adherence to legal traditions and frameworks. Moreover, participants recognized the potential of multi-turn interactions to help users better articulate their legal issues and access relevant information. The following sections present a detailed analysis of these qualitative findings, organized around the central themes that emerged from the workshop data.
@@
  • Refrain from making definitive judgments about the legal consequences of a specific case. Providing relevant laws is fine (e.g., driving under influence (DUI) is illegal in Washington) but applying it to specific user situations constitutes opinion (e.g., falling asleep in the driver’s seat in the parking lot after drinking alcohol could be a DUI) (P2, P13, P17, P19).
@@
In essence, legal opinion encompasses interpretive, judgment-driven analysis that is often value-laden and forward-looking, whereas legal information involves reporting objective laws and past rulings without subjective assessment. To understand this distinction, we can draw upon the widely-used legal analysis tool known as IRAC (Issue-Rule-Analysis-Conclusion) (Metzler, 2002). IRAC entails (1) identifying the legal issue, (2) stating the rule that applies, (3) analyzing how the particular facts interact with the stipulations of the rule, and (4) finally deducing the conclusion (Center, 2001). Our findings suggest that LLM systems focusing on issue and rule identification provide fact-finding “information,” while analysis and conclusions may cross into tailored “opinion,” as illustrated in Figure 5. However, it is important to acknowledge the complexity of distinguishing between legal information and opinion, as the line between the two can often be blurred in practice, as exemplified by cases such as Grievance Comm. of Bar v. Dacey (boo, 1966), where the court found that publishing a booklet providing trust and tax information crossed the line into unauthorized legal opinion. This demonstrates that the distinction between legal information and opinion is not always clear-cut, and careful consideration must be given to the specific context and the information provided by LLM systems.
@@
While cautioning against detailed legal opinions, participants suggested that LLMs could offer a better user experience compared to traditional search engines. P20 noted that users would not find it helpful if LLM systems “vomit a whole lot of knowledge.” The most promising and heavily-discussed possibility during the workshops is leveraging multi-turn interactions, allowing LLMs to ask follow-up questions and clarify users’ legally meaningful inquiries. This idea emerged as participants expressed frustration with missing case facts: “I don’t think there’s enough information to go off of, and that depending on the details that come out, it could change the analysis, therefore the outcomes.” (P13) Participants emphasized that legal contexts are inherently complex (P11), and lawyers often spend considerable time eliciting relevant facts and identifying the “right questions to ask” (P12). They felt that LLM-mediated dialogues could streamline time-consuming processes such as “screening interviews” (P12), “first calls” (P14), or “intake meetings” (P15). By engaging in multi-turn interactions, LLMs could help users refine their questions, focus on key aspects of their cases, and seek relevant expertise. However, some warned that LLM developers should exercise caution when eliciting extensive personal information from users, given confidentiality concerns (P13, P16). While identifying legal issues and relevant rules likely falls within the realm of permissible legal information, the line between information and opinion remains blurred. P16 argued that narrowing down factual patterns and applying rules engages deep judgement, stating “You’re starting to make the AI become your lawyer.”
@@
Experts suggested several principles for providing LLM legal guidance. Some principles directly align with emerging literature on transparency (Kim et al., 2023b), user satisfaction (Kim et al., 2023a), and cautions about anthropomorphism (Shanahan et al., 2023; Ma et al., 2024). The principles outlined below represent the most prominent and frequently discussed ideas that emerged from the expert discussions.
  • Don’t Pretend to Be Human: LLM systems should not behave like a human and cause misrepresentations, as that can create issues around transparency, over-reliance, and managing user expectations.
  • Caveat Constraints: LLM systems should provide various caveats on its limitations, such as that its capabilities are constrained, the conversation is not privileged, and it is working off of incomplete information.
  • Avoid Potential Harm: LLM systems should refrain from providing recommendations that could potentially cause harm to users or others. This includes avoiding guidance that may lead to harmful real-world actions, as well as minimizing the risk of emotional or psychological harm that could result from the system’s responses.
  • Respect the Justice System: LLM systems should not give answers that could enable users to intentionally avoid law enforcement or oversight.
  • Avoid Unethical Answers: LLM systems should not make any outputs that could promote dishonesty, deception, fraud, impersonation, or other unethical behaviors that could get users into trouble.
  • Be Transparent: LLM systems should be able to explain the outcome it generated and point to the specific areas of datasets it relied on.
  • Avoid Appearance of Impropriety: The appearance of impropriety refers to a situation that may appear corrupt or unethical to an impartial observer. For example, the LLM system should not endorse or promote its creators, the AI companies involved in its development, or any other entities that could be perceived as influencing the system’s outputs. The LLM responses should be objective and impartial, focusing solely on providing accurate and helpful information to users.
@@ 4.3. Summary of Results
Our analysis uncovered 25 distinct dimensions to ensure safe and effective LLM legal advice, spanning four key categories: (1) User attributes and behaviors, (2) Query characteristics, (3) AI capabilities, and (4) Social impacts. Experts deliberated with each other and through points of consensus to produce this rich set of considerations. However, experts expressed limited consensus on how LLM systems should actually respond, given these nuanced factors. Some remained resistant to any LLM involvement in legal questions, while others envisioned more helpful LLM assistance models that increase access to information. Most debates surrounded distinguishing information versus opinion, and the majority felt that providing factual legal information is appropriate. Some participants suggested using LLMs’ conversational capabilities to help users refine questions and identify relevant laws through follow-up questions, similar to initial consultations with attorneys.
@@ 5. Discussion
Constructing LLM policies does not exist in a technocratic silo. Rather, it demands a cross-disciplinary approach that synthesizes insights from domain experts. Our research demonstrates that engaging legal experts in case-based deliberation is an effective method for translating professional knowledge and clinical experience into a concrete set of considerations for LLM policies. Our 4-dimension framework we have developed provides a useful analytical lens that can be applied to future exploring LLM policies in the legal advice and other professional contexts. Through this approach, we argue that policymakers can derive valuable insights to inform LLM policies grounded in the centuries-old wisdom and experience of the legal profession, while also accounting for the challenges presented by LLM technologies.
5.1. Benefits of Case-based Deliberation Methods
@@
Our research process underscores several advantages of grounded case deliberation for eliciting expert considerations. Preparing realistic scenarios, while laborious, proved invaluable in quickly engaging experts with targeted queries related to their clinical experience. The cases allowed experts to examine granular concerns around singular situations as well as overarching technical and legal constraints, producing a more concrete set of contextual factors for AI developers, beyond theoretical and high-level principles in prior research (Solaiman et al., 2023; Antoniak et al., 2023; Ganguli et al., 2022). Finally, the collective deliberation itself revealed critical hidden dimensions and elicited justifications that shed new light on existing dimensions. As experts built on each other’s points, they realized overlooked issues or limitations in their own initial analyses. This interplay sharpened considerations and revealed nuances around balancing risks and benefits in varied situations. The combination of realistic cases and collaborative discourse resulted in more fine-grained, practice-informed insights compared to de-contextualized surveys or high-level principles.
5.2. Charting Novel Legal Considerations
One of our contributions is to shed light on existing legal and ethical barriers to LLMs’ legal advice which have been overlooked in the literature. Section 4.1.3 reveals that users lack confidentiality and accountability protections governing attorney advice: Conversations with AI systems risk disclosure in legal proceedings and inaccurate guidance evades professional negligence liability. Moreover, as Section 4.2.2 explains, UPL regulations prohibit non-lawyers from advising in many states, carrying criminal penalties. To circumvent the current legal risks , one could imagine LLM systems designed like private counsels advising single parties, rather than serving all users uniformly like ChatGPT. In such case, LLM systems could come to resemble proprietary services, with corresponding confidentiality and liability assurances. However, the legal conservatism may change in the future as UPL rules have already faced criticism for limiting affordable access to legal help (Denckla, 1998; Rotenberg, 2012; Association, 1995). The EU AI Act’s categorization of AI legal assistance tools as “high-risk,” which subjects them to heightened responsibilities instead of banning them outright, may speak to this potential shift (European Commission, 2021).
5.3. Learning from Time-tested Wisdom
@@
Leveraging accumulated expertise in professional communities can help sidestep painful mistakes (Antoniak et al., 2023). In our research, UPL does not only constrain the possibility of LLM legal advice but also provides long-standing distinction criteria between information versus opinion, as merely providing legal information has not been historically punished as a UPL violation (cal, 1970; Denckla, 1998; cal, 2007). For example, the Texas Court provides guidelines for court staff and illustrative examples like in Table 2. These examples show subtle differences between information and opinion, which resembles the red-teaming approach to distinguish harmful user prompts (Ganguli et al., 2022; Achiam et al., 2023). Furthermore, legal scholars have explored legally justifiable AI advice under UPL, attorney-client privilege, and other doctrines (Haupt, 2019; Wendel, 2019; Spahn, 2017). Wendel stated that the “core lawyering functions” such as recommending the course of action or drafting contracts cannot be delegated to AI agents due to technical limitations and accountability deficits (Wendel, 2019). This demonstrates how principles accumulated over centuries of legal scholarship now inform responsible LLM systems and the call for cross-disciplinary collaborations.
@@
Procedure Can you tell me how to file a small claims action? Can you tell me whether it would be better to file a small claims action or a civil action?
Definition What does “certificate of service” mean? My neighbors leave their kids at home all day without supervision. Isn’t that child neglect?
Forms I need to file for divorce and I have no idea where to begin. Is there some place I can go to find out how to get started? The self-help divorce petition says I should list any gifts as my separate property. Should I list the money that my parents gave me last month as my separate property?
Options What can I do if I cannot afford to pay the filing fee? My ex-husband hasn’t paid the debts that he agreed to pay in our divorce settlement. Can I be made responsible for this debt?
Table 2. Examples of impermissible questions that requires legal opinion (of Court, 2015). Remarkably similar to how red-teaming in LLM development identifies harmful user inputs (Ganguli et al., 2022; Achiam et al., 2023), this edited list (compiled from Texas law clerk resources) distinguishes between permissible and forbidden questions Texas court personnel can answer.
5.4. Applicability to Other Professional Domains
While each possessing unique dimensions, domains like medicine, mental health, law, and finance share common threads around high-stakes real-world impact and historical reliance on licensed specialists for advice. We believe that our research methods and 4-dimension framework give illustrative guidance to further research in other professional domains. As this research demonstrates how case-based deliberation methods can unravel complex professional ethics, researchers could adopt similar processes engaging mental health counselors, financial advisors, or medical professionals. Tapping into the clinical experience and integrity of practitioners through structured deliberation based on realistic cases can help produce tailored dimensions and guidelines for responsible LLM advice respective to each profession. Building upon this foundation, our 4-dimension framework—(1) user, (2) query, (3) AI capabilities, and (4) impact—could be adapted and applied across various professional domains. The (1) user, (2) AI, and (4) impact dimensions can be applied in other domains with minimal modifications. However, the query dimension requires more customization to address the typical requests of clients, terminology, and satisfactory responses in each field.
5.5. Limitations and Future Research
Our study has several limitations. First, our expert sample predominantly focused on practitioners familiar with the US legal system. Ethical considerations around appropriate AI assistance may differ across different legal systems and cultures. Second, our participants’ responses are conditioned by their prior experience with state-of-the-art LLM technology, such as ChatGPT empowered by GPT-4. Experts’ evaluations of the appropriateness of LLM legal advice may evolve in the future, based on technological innovations, which could be an avenue for future research. Third, we did not engage end-users like clients of legal services. Future work can specifically investigate end-user perceptions to compare and contrast with our expert-informed results. Finally, while our taxonomy conceptualizes a concrete set of dimensions, how these dimensions could change the appropriateness of LLM responses remains unexplained. This may require larger-scale empirical analysis on public assessments across diverse pairings of cases and responses.
@@ 6. Conclusion
Today, LLM chatbots are increasingly capable of providing users with advice in a wide range of professional domains, including legal advice. However, what constitutes an appropriate LLM-generated response to legal queries, where both required expertise and resulting consequences are high? To explore this, we conducted workshops with 20 legal experts using methods inspired by case-based reasoning to encourage deliberations around appropriate LLM responses to legal queries in practice. Our contributions are threefold. First, we presented a set of 25 key dimensions, synthesized from expert deliberations, that impacted LLM response appropriateness in the legal domain. Second, we shared experts’ recommendations for LLM response strategies and guiding principles for generating appropriate responses—these centered around helping users identify and prepare salient information for legal proceedings rather than recommending specific legal actions. Finally, we posit that our case-based method has utility in engaging expert perspectives on LLM response appropriateness in professional domains beyond the legal sphere. Taken together, our work sets an empirical foundation for translating domain-specific professional knowledge and practices into policies to steer real-world LLM behavior in a more responsible direction.